OpenStudy (anonymous):

Check it please: Find the (shortest) distance from the origin to the parabola y=4-x^2

OpenStudy (anonymous):

I got z= sqrt( (x-0)^2 + (y-0)^2 )

OpenStudy (anonymous):

Then z= sqrt( (x-0)^2 + (4-x^2-0)^2 )

OpenStudy (anonymous):

Then z= sqrt( (x)^2 + (4-x^2)^2 )

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

Find the interecepts and then take the square root of the sum of their squares

jimthompson5910 (jim_thompson5910):

The next step would be to derive z with respect to x Then you find the roots of the derivative function you found

OpenStudy (anonymous):

dz/dx = (x(2x^2-7))/(sqrt(x^4-7x^2+16)) ??

jimthompson5910 (jim_thompson5910):

close, but not quite

jimthompson5910 (jim_thompson5910):

oh wait, i see how you got that

jimthompson5910 (jim_thompson5910):

yes you are correct

jimthompson5910 (jim_thompson5910):

now set this equal to zero and solve for x

OpenStudy (anonymous):

so x=sqrt(7/2)

jimthompson5910 (jim_thompson5910):

x = 0 is also a root

jimthompson5910 (jim_thompson5910):

those are your critical values now you have to use the first or second derivative test to see which one of those critical values are minimums

OpenStudy (anonymous):

the sqrt(7/2) is

jimthompson5910 (jim_thompson5910):

and so is -sqrt(7/2)

jimthompson5910 (jim_thompson5910):

and this makes sense because you have symmetry with respect to the y-axis

OpenStudy (anonymous):

we plug x into sqrt(x^4-7x^2+16)

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

to find the distance

OpenStudy (anonymous):

and get + or - 1.936494154 ?

jimthompson5910 (jim_thompson5910):

I'm getting the same

OpenStudy (anonymous):

Woot thanks a bunch for the help!

jimthompson5910 (jim_thompson5910):

you're welcome